Consider the following forward contracts: Contract A: Long AUD against USD, notional amount AUD 10,000, forward exchange rate 0.80 USD per AUD. Contract B: Short AUD against USD, notional amount AUD 50,000, forward exchange rate 0.75 USD per AUD. Suppose both contracts are maturing today, and the spot exchange rate is 0.70 USD per AUD. Please calculate the profit or loss (P&L) on each contract. P&L for Contract A - -1000, or 1000 or none P&L for Contract B - -2500, 2500 or none

Now suppose you can enter either Contract A or Contract B today with a three-month maturity. In three months, you believe the spot exchange rate will be 0.82 USD per AUD. Which contract do you prefer? A OR B
Answer #1

Contract A: Long AUD against USD

Forward exchange Rate=0.8USD per AUD

Cost of buying 10000 AUD in forward =0.8*10000 USD=8000 USD

Spot exchange rate =0.70USD per AUD

Cost of buying 1000 AUD in spot=0.70*10000=7000 USD

Profit/(Loss)=7000-8000=-1000 USD

ANSWER: Contract A: -1000 USD

Loss of 1000 USD

Contract B: Short AUD against USD

Forward exchange Rate=0.75USD per AUD

Cost of 50000 AUD in forward =0.75*50000 USD=37500 USD

Spot exchange rate =0.70USD per AUD

Cost of 50000 AUD in spot=0.70*50000=35000 USD

Profit/(Loss)=37500-35000=2500 USD

ANSWER: Contract B: 2500 USD

Profit of 2500USD

If spot rate after 3 months is0.82 USD per AUD

Contract A : Forward Rate =0.8 USD per AUD

Forward rate for BUYING is Less than Spot rate

There will be Profit in Buying in Forward

Profit =10000*(0.82-0.8)=200 USD

Contract B: Forward Rate 0.75 USD per AUD

Forward Rate is less than Spot rate

There will be Loss in Selling in Forward

Hence CONTRACT A should be preferred